OpenStudy (anonymous):

Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0

OpenStudy (anonymous):

x^3 or x3?

OpenStudy (anonymous):

-x^3 + 5x^2 - 11x + 55 = 0

OpenStudy (ddcamp):

Here we should factor by grouping. \[0=[-x^3+5x^2]+[-11x+55]\] Pulling out the common factors on both sides, we're left with: \[0=x^2(-x+5) + 11x(-x+5)\] Which gives us: \[0=(x^2+11x)((5-x)\] Pulling an x out of the left term, we get \[0= x(x+11)(5-x) \] Now find the zeroes by setting each factor equal to zero

OpenStudy (anonymous):

I have no idea how to do that...

OpenStudy (ddcamp):

Honestly, factoring by grouping is something that they just put in the book for no good reason. It only works on specific equations, and you won't be using it much later on.

OpenStudy (anonymous):

Oh... that's... tragic >-< but that's a relief lol thank you